### v1

# Distributed implementation of tensor-product selected configuration interaction

Abstract
In recent years, the hybrid “QC+HPC” strategy—where quantum computers screen important determinants, followed by exact diagonalization on classical computers—has shown great potential in the study of strongly correlated systems in quantum chemistry. Last year, an IBM team proposed a novel scheme that utilizes quantum computers to select important bit strings that are then used to construct a spin-adapted determinant space via tensor products. Inspired by this, we have specifically designed a completely new algorithm for this tensor-product selected configuration interaction (SCI). Moreover, for the first time worldwide, we have implemented distributed storage of the CI vector in an SCI program, enabling efficient handling of large-scale computation. Since this study is independent and does not involve determinant selection by quantum computers, we employed our SCI program to conduct full configuration interaction (FCI) computations. Our FCI calculations for N2 (aug-cc-pVDZ) under D2h symmetry and CN (cc-pVTZ) under C2v symmetry, involving $1.47\times 10^{11}$ and $4.86\times 10^{11}$ determinants, respectively, exceed the previous record of $2\times 10^{9}$ determinants computed with the DICE program [J. Chem. Phys. 149, 214110 (2018)] by more than two orders of magnitude. These results set a new benchmark for SCI computations and lay the groundwork for further advancements in SCI methods.
1 Introduction
The Full Configuration Interaction (FCI) is an exact method for solving the Schrödinger equation in quantum chemistry. 1, 2, 3, 4 However, due to its exponential scaling with system size, FCI remains limited to small molecules with small basis sets. By the use of a distributed CI vector, the largest system calculated to date is propane with the STO-3G basis set, involving up to $1.31\times 10^{13}$ determinants 4.
In practical applications, the coupled cluster singles and doubles with perturbative triples [CCSD(T)] method 5 is known as the gold standard for weakly correlated systems owing to its balanced accuracy and efficiency. While achieving accurate calculations for strongly correlated systems remains the most formidable challenge in quantum chemistry, the selected configuration interaction (SCI) method has long been employed and is recognized as one of the most promising methods for addressing this issue. As early as 1973, Huron et al. introduced the Configuration Interaction using a Perturbative Selection made Iteratively (CIPSI) method 6 in which candidate determinants are selected by rigorously estimating their energy contributions via second-order perturbation theory. Although this selection scheme is considered reliable, its high computational cost precludes its application to larger molecules. Subsequently, methods such as Adaptive Sampling CI (ASCI) 7, 8, and Semistochastic Heat-Bath CI (SHCI) 9, 10, 11 were developed, which can essentially be regarded as approximations to CIPSI.
In recent years, with the rise of quantum computing (QC), several research groups have attempted to rapidly screen important determinants using quantum devices. However, the noise inherent in quantum computers can lead to the loss of some determinants, preventing the generation of wavefunctions that strictly correspond to the eigenstates of the $\hat{S}^{2}$ operator. To address this issue, last year a team from IBM proposed an innovative strategy: using quantum computers to select important bit strings, and then constructing the complete determinants via tensor products. 12 In this paper, we refer to this variant of SCI,which is characterized by this structure, as Tensor-Product SCI (TPSCI).
Since CI calculations are variational, including more determinants generally improves accuracy; however, the associated memory and computational costs for diagonalization increase dramatically. Currently, the widely used DICE program, 9, 10 which distributes the nonzero Hamiltonian matrix elements across nodes, enables calculations with up to $2\times 10^{9}$ determinants. Nevertheless, DICE employs replicated storage for the CI vector, so that the memory available on a single node becomes a bottleneck, thereby limiting the maximum feasible CI vector size despite multi-node execution.
In this work, we have, for the first time worldwide, implemented distributed storage of the CI vector in an SCI program, overcoming single-node memory limitation and enabling the use of the entire supercomputer’s available memory for larger-scale calculations. Moreover, we have designed a completely new, highly efficient algorithm specifically for TPSCI, leveraging its unique tensor-product structure, and further optimized the program for efficient parallel execution on the supercomputer Fugaku. We also validate the use of single precision to reduce computational cost and memory use; this is a strategy pursued for Configuration Interaction as early as the work of Pakiari and Handy, 13 and in line with current trends in correlated methods. 14 The combination of low memory footprint and low precision operations make our work promising for application to leadership class supercomputers based on GPUs.
Ultimately, we computed the FCI energies of N2 (aug-cc-pVDZ) and CN (cc-pVTZ), which, under D2h and C2v symmetry, respectively, involve $1.47\times 10^{11}$ and $4.86\times 10^{11}$ determinants. These results significantly exceed the previously reported maximum SCI scale of $2\times 10^{9}$ determinants computed with the DICE program. 11 These achievements lay a solid foundation for the future development and application of TPSCI. Meanwhile, as previous FCI calculations for N2 15 and CN 16 were performed with the smaller cc-pVDZ basis set, our results serve as valuable benchmark data for further research.
2 CI Vector Distributed Algorithm for TPSCI
In this section, we will describe the parallel algorithm we have designed for TPSCI, including the important optimizations required for large-scale application.
2.1 Distributed CI Vector Storage
In the framework proposed by IBM, only important bit strings are selected, which means that the chosen $\alpha$ and $\beta$ bit strings are identical. The determinant space constructed via tensor products can thus be used to generate closed-shell singlet wavefunctions that strictly correspond to the eigenstates of the $\hat{S}^{2}$ operator.
In our program, the selected $\alpha$ and $\beta$ bit strings are allowed to differ. Suppose their sets are defined as $\{S_{w}^{\alpha}\mid w=1,2,\dots,x\}$ and $\{S_{u}^{\beta}\mid u=1,2,\dots,y\}$, where $S^{\alpha}$ and $S^{\beta}$ are the sets of selected $\alpha$-spin and $\beta$-spin bit strings, respectively, with elements denoted as $S_{w}^{\alpha}$ and $S_{u}^{\beta}$. The determinant space constructed via tensor products is then given by:
$\{|S_{w}^{\alpha}\rangle\otimes|S_{u}^{\beta}\rangle\mid w=1,\dots,x;\,u=1,\dots,y\}.$ (1)
The ordering of $\{S^{\alpha}\}$ and $\{S^{\beta}\}$ can be freely chosen. However, once determined, the index order of the determinants in the determinant space is also fixed. Specifically, the $K$-th determinant is denoted as $|D_{K}\rangle$ and given by:
$|D_{K}\rangle=|S_{w}^{\alpha}\rangle\otimes|S_{u}^{\beta}\rangle,\quad w=\frac{K}{y}+1,\quad u=\text{MOD}(K,y).$ (2)
When the number of determinants exceeds the storage capacity of a single node, the CI vector must be distributed across nodes. In this work, the term CI vector refers to the array of CI coefficients stored in the same order as the determinant indices. As illustrated in Fig. 1, we distribute the determinants based on their $\alpha$ bit strings while ensuring an approximately even allocation, a strategy known as memory balance.
Since the determinants are indexed such that $\alpha$ bit strings vary first, followed by $\beta$ bit strings, we define a segment as the subset of the CI vector corresponding to all determinants sharing the same $\alpha$ bit string. In this work, we denote the total number of processes (each node runs exactly one process) as $N$. Let $m_{p}$ denote the number of $\alpha$ bit strings assigned to process $p$ (where $p=0,1,\dots,N-1$), with $\sum^{N-1}_{p=0}m_{p}=x$. Then, each process contains $m_{p}$ segments and thus is responsible for a total of $m_{p}\times y$ determinants. As each determinant is fully defined by its index through Eq. (2), our program does not require additional memory to store determinant composition information.
Figure 1: Distribution of the CI vector into segments across $N$ processes. Each process holds a strict subset of $m_{p}$ of the $\alpha$ strings. On a local process, the determinants are divided into segments (represented by vertical arrows), where a segment is composed of all determinants that share the same $\alpha$ string. For this figure, we define a displacement index $d_{p}=\sum_{i=0}^{p-1}m_{i}$ to show the offset of $\alpha$ strings on a given process.
2.2 Distributed Matrix-Vector Multiplication
In this work, we employ the traditional Davidson diagonalization method 17 for CI calculations, where the core computational step is evaluating $\mathbf{W}=H\cdot\mathbf{U}$, with Hamiltonian matrix $H$, and CI vectors $\mathbf{W}$ and $\mathbf{U}$.
The most advanced SCI implementations to date follow two main strategies: some store the nonzero Hamiltonian elements in a distributed manner across nodes (e.g., DICE), while others compute the Hamiltonian matrix elements on-the-fly (e.g., QUANTUM 18). However, both approaches store the CI vector entirely on a single node without exception. The advantage of these approaches is that each computing node performs a simple matrix-vector multiplication; at the end of each Davidson iteration, the effective Hamiltonian matrix is efficiently constructed using an MPI_ALLREDUCE operation. However, a key limitation is that the CI vector size is strictly constrained by the memory of a single node, posing a major bottleneck in utilizing supercomputing resources for larger determinant spaces.
In our program, the CI vector is distributed across nodes, and each node runs exactly one process. In this setting, each node $p$ only stores its corresponding CI vector segments. The computation of $\mathbf{W_{p}}$ is then given by:
$\mathbf{W_{p}}=\sum_{q}H\cdot\mathbf{U_{q}}.$ (3)
This implies that, in a single Davidson iteration, node $p$ theoretically needs to sequentially fetch data from all other nodes to compute $\mathbf{W_{p}}$. We define each individual fetch-and-compute operation (i.e., computing $\mathbf{W_{p}}\mathrel{+}=H\cdot\mathbf{U_{q}}$) as one step. Since there are $N$ computing nodes, each Davidson iteration consists of $N$ steps. Unlike the implementation in QUANTUM, we precompute and store the diagonal elements of the Hamiltonian matrix in memory, while the off-diagonal elements are computed on-the-fly during the Davidson iterations.
During the $N$ computation steps within a single Davidson iteration, each node operates independently without global synchronization. For example, once node $p$ completes the computation $\mathbf{W}_{p}\mathrel{+}=H\cdot\mathbf{U_{q}}$, it immediately fetches $\mathbf{U_{r}}$ from node $r$ and proceeds with $\mathbf{W_{p}}\mathrel{+}=H\cdot\mathbf{U_{r}}$. This design is crucial because computation times vary across different steps and nodes. Enforcing synchronization after each step would lead to unnecessary waiting and degrade computational efficiency. For data communication, we adopt the MPI_GET operation, a type of one-sided communication, allowing nodes to fetch data asynchronously without interrupting the computation on the target node, further improving efficiency.
2.3 Efficient Hamiltonian Computation for TPSCI
Due to the Slater-Condon rules, $H$ is a highly sparse matrix, so minimizing operations that involve zero elements is crucial for efficiency. In the DICE program, the nonzero Hamiltonian elements are precomputed and stored in a distributed manner along with their corresponding indices, allowing direct computations with these elements and avoiding operations on zeros.
While such precomputation strategies are effective, they require significant memory to store the nonzero elements. In contrast, our approach avoids explicit storage of Hamiltonian elements and instead exploits the inherent sparsity of $H$ to eliminate unnecessary computations. Here, we first analyze the structure of Hamiltonian matrix elements in SCI, which provides the foundation for our efficient computation strategy.
The Hamiltonian matrix element $\langle D_{w}|H|D_{u}\rangle$ can be expressed as $\langle S_{w}^{\alpha}|\langle S_{w}^{\beta}|H|S_{u}^{\alpha}\rangle|S_{u}^{\beta}\rangle$. According to the Slater-Condon rules, this matrix element can be nonzero only if:
$\text{DIFF}(S_{w}^{\alpha},S_{u}^{\alpha})+\text{DIFF}(S_{w}^{\beta},S_{u}^{\beta})\leq 2,$ (4)
where $\text{DIFF}(S_{w},S_{u})$ denotes the number of electron differences between bit strings $S_{w}$ and $S_{u}$. To classify these terms concisely, we introduce the notation $[a,b]$, where $a=\text{DIFF}(S_{w}^{\alpha},S_{u}^{\alpha})$ and $b=\text{DIFF}(S_{w}^{\beta},S_{u}^{\beta})$. For example, the term [1,1] corresponds to $\text{DIFF}(S_{w}^{\alpha},S_{u}^{\alpha})=1$ and $\text{DIFF}(S_{w}^{\beta},S_{u}^{\beta})=1$. Throughout this work, we use the shorthand notations [2,0], [1,1], [1,0], [0,2], [0,1], and [0,0] to represent a total of six terms in which the matrix elements may be nonzero. The first five terms correspond to off-diagonal elements of the Hamiltonian matrix, which are always computed on-the-fly. The last term, [0,0], corresponds to diagonal elements, which, although also computable on-the-fly, are precomputed and stored in memory in our program to improve efficiency.
Data: p the current process, q the current step.
            SINGLE_LINK and DOUBLE_LINK lists.
1 for $S_{w}^{\alpha}$leftarrow $S^{\alpha}_{p,1}$ to $S^{\alpha}_{p,m_{p}}$ do
2      for $S_{u}^{\alpha}$leftarrow $S^{\alpha}_{q,1}$ to $S^{\alpha}_{q,m_{q}}$ do
3           case $\leftarrow$ DIFF($S_{w}^{\alpha}$, $S_{u}^{\alpha}$)
4           if case = 2 then
5                for $S_{w}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{y}$ do
6                     $S_{u}^{\beta}\leftarrow S_{w}^{\beta}$
7                     Calculate [2,0] term
8
9                     end for
10
11                     else if case=1 then
12                          for $S_{w}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{y}$ do
13                               foreach $S_{u}^{\beta}$in SINGLE_LINK(:,$S_{w}^{\beta}$do
14                                    Calculate [1,1] term
15
16                                    end foreach
17                                   $S_{u}^{\beta}\leftarrow S_{w}^{\beta}$
18                                    Calculate [1,0] term
19
20                                    end for
21
22                                    else if case=0 then
23                                         for $S_{w}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{y}$ do
24                                              foreach $S_{u}^{\beta}$in SINGLE_LINK(:,$S_{w}^{\beta}$do
25                                                   Calculate [0,1] term
26
27                                                   end foreach
28                                                  $S_{u}^{\beta}\leftarrow S_{w}^{\beta}$
29                                                   Calculate [0,0] term
30
31                                                   end for
32
33                                                   end if
34
35                                                   end for
36                                                  $S_{u}^{\alpha}$leftarrow $S_{w}^{\alpha}$
37                                                   for $S_{w}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{y}$ do
38                                                        foreach $S_{u}^{\beta}$in the q-th batch of DOUBLE_LINK(:,$S_{w}^{\beta}$do
39                                                             Calculate [0,2] term
40
41                                                             end foreach
42
43                                                             end for
44
45                                                             end for
Algorithm 1 Computation algorithm for TPSCI on local process $p$ at step $q$. This algorithm describes how to compute $\langle D_{w}|H|D_{u}\rangle$, where $|D_{w}\rangle$ and $|D_{u}\rangle$ are determinants from processes $p$ and $q$, respectively.
Our novel algorithm for TPSCI is shown in Algorithm 1. Each Davidson iteration consists of $N$ steps. In step $q$, the local process $p$ performs the computation using data fetched from process $q$, which is detailed in Algorithm 1.
To eliminate unnecessary computations, we introduce the SINGLE_LINK and DOUBLE_LINK arrays, which store precomputed single and double excitation connections among $\{S_{w}^{\beta}\mid u=1,2,\dots,y\}$. These arrays ensure that Hamiltonian evaluations only occur when the electron difference between $|D_{w}\rangle$ and $|D_{u}\rangle$ does not exceed 2, directly eliminating invalid cases and improving efficiency. This optimization is enabled due to the unique structure of TPSCI: for each $\alpha$ bit string, the associated $\beta$ bit string set remains unchanged across determinants, allowing excitation connections to be precomputed and reused efficiently.
As the DOUBLE_LINK array can be extremely large, it is distributed nearly evenly across all processes based on its indices. To keep track of this distribution, an auxiliary array is introduced to record the starting and ending indices of DOUBLE_LINK assigned to each process. Consequently, the computation of the [0,2] term is performed batch by batch over the $N$ steps (as shown in Lines 27-32 of Algorithm 1). Since the calculation of the [0,2] term is performed only in the loop over $S_{w}^{\alpha}$, each process $p$ will fetch data from the target process $q$ for the DOUBLE_LINK only once per Davidson step.
Roughly speaking, the computational cost of TPSCI scales as $N_{\mathrm{SCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}$, where $N_{\mathrm{SCI}}$ is the number of selected determinants, $N_{\mathrm{occ}}$ the number of occupied orbitals, and $N_{\mathrm{vir}}$ the number of virtual orbitals. This can be further reduced by considering the sparsity of excitation links. The sparsity of single excitation links roughly scales as $\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}$, in which $N_{\mathrm{FCI}}$ is the FCI dimension, and thus the overall scaling is more precisely given by $N_{\mathrm{SCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot N_{\mathrm{SCI}}/N_{\mathrm{FCI}}$.
2.4 Strategies for Efficient Parallel Execution in TPSCI
As described above, we distribute the CI vector across nodes to overcome single-node memory limitations, enabling computations to fully utilize the total available memory. However, this distributed storage introduces new challenges: a computation that could previously be completed in a single step now requires $N$ steps, with each step requiring data retrieval from a target process [as shown in Eq. (3)]. As the number of nodes increases, the overall MPI communication volume seemingly scales as $N_{\text{SCI}}^{2}$, making communication latency a potential bottleneck for computational efficiency. To address this issue, we have developed a series of strategies to reduce communication overhead and improve parallel execution.
2.4.1 Avoiding Unnecessary MPI Data Transfers
In our previous analysis, we mentioned that each local process $p$ needs to fetch data from other processes. However, not all data from the target process is required. Specifically, if all $\alpha$ bit strings assigned to process $p$, namely $\{S^{\alpha}_{p,1},S^{\alpha}_{p,2},\dots,S^{\alpha}_{p,m_{p}}\}$, satisfy $\text{DIFF}(S^{\alpha}_{w},S^{\alpha}_{u})>2$ for any $S^{\alpha}_{w}$ in this set, then process $p$ does not need the segment corresponding to $S^{\alpha}_{u}$.
In our implementation, such information is precomputed and stored, allowing each process to identify and skip unnecessary data transfers, thereby significantly improving communication efficiency. As a result, the MPI communication volume for process $p$ roughly scales as $m_{p}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}\cdot y$. It can be easily shown that the computational cost and the data communication volume on each process share the same scaling. Consequently, the overall MPI communication volume across all nodes scales as $N_{\mathrm{SCI}}$ rather than $N_{\mathrm{SCI}}^{2}$, providing a solid foundation for the efficient parallel execution of TPSCI on supercomputers.
2.4.2 Minimizing Long-Distance Communication in MPI
Each node in the supercomputer is assigned a unique node identity number (ID) ranging from 00 to $N-1$. When suitably ordered, messages sent between nodes with closer IDs will on average require fewer hops compared to those with larger ID differences.
Based on this, we sort the $\alpha$ bit strings according to their excitation levels relative to the Hartree-Fock (HF) $\alpha$ bit string, and then assign them to nodes in increasing order. Hence, lower-excitation $\alpha$ bit strings are assigned to lower-numbered nodes, whereas higher-excitation strings are assigned to higher-numbered nodes.
As a result, communication between distant nodes is significantly reduced because the excitation differences between their assigned $\alpha$ bit strings are more likely to be all greater than 2, making communication unnecessary. This strategy is crucial for reducing MPI communication congestion, a well-known challenge in large-scale high-performance computing, by ensuring that each node primarily fetches data from nearby nodes, thereby improving parallel efficiency.
2.4.3 Achieving Computational Load Balance in MPI
In MPI-based parallel computing, maintaining computational load balance is crucial because the slowest process determines the overall execution time. The challenge is how to estimate each process’s workload in advance. As shown in Algorithm 1, the most time-consuming computation is the evaluation of the [1,1] term. Therefore, the number of single excitations associated with each $\alpha$ bit string can be used as a rough estimate of the computational workload for its corresponding segment. Consequently, the total computational workload of a process can be estimated by summing the single excitation counts of all $\alpha$ bit strings assigned to it.
As described above, we assign the $\alpha$ bit strings to processes in increasing order of their excitation levels relative to the HF $\alpha$ bit string. This ordering is crucial in our TPSCI program to minimize long-distance communication between nodes. However, balancing memory usage and computational workload simultaneously presents a challenge. Since the total number of $\alpha$ bit strings is fixed, the average number per node is also predetermined. Similarly, the total number of single excitation counts of all $\alpha$ bit strings is fixed, meaning that the average single excitation count per process is also predetermined.
To quantify the balance between memory and computation, we define two expansion factors: the memory expansion factor, which is the ratio of the assigned $\alpha$ bit string count to the average, and the computation expansion factor, which is the ratio of the assigned single excitation counts to the average. Enforcing strict memory balance (i.e., keeping the memory expansion factor close to 1) can lead to severe computational imbalance, causing some nodes to finish significantly earlier and remain idle while waiting for others. Conversely, enforcing strict computational balance (i.e., keeping the computation expansion factor close to 1) may result in a highly uneven memory distribution, with some nodes exceeding their memory limits, forcing an increase in the total number of nodes and leading to inefficient resource usage.
To address this issue, we adopt a balanced strategy where both expansion factors are kept moderately greater than 1, ensuring that neither memory imbalance nor computational imbalance becomes excessive.
2.4.4 Overlapping Computation and Data Transfer to Reduce MPI Latency
As mentioned above, each node in our program runs a single process. Modern computers are typically multi-core; for example, each node on supercomputer Fugaku has 48 CPU cores. Even though we minimize unnecessary data transfers and reduce long-distance communication, some degree of MPI communication congestion remains unavoidable. The question then arises: how can we further mitigate its impact on the overall computation wall time?
To mitigate this issue, one core is assigned to perform MPI data transfers, fetching the required data for the next step, while the remaining 47 cores perform parallel computation for the current step using OpenMP. Let $t_{\text{cal}}$ and $t_{\text{data}}$ represent the computation time and the data retrieval time of one step, respectively. If $t_{\text{data}}>t_{\text{cal}}$, we define the delay time to the wall time as $t_{\text{delay}}=t_{\text{data}}-t_{\text{cal}}$. Conversely, if $t_{\text{data}}\leq t_{\text{cal}}$, then $t_{\text{delay}}=0$, since MPI communication congestion during data transfer—if it occurs—does not prolong the wall time. The total delay time for a Davidson iteration, denoted as $T_{\text{delay}}$, is given by the sum of $t_{\text{delay}}$ over all $N$ steps.
2.4.5 “Check if Busy” Strategy for Communication
During data fetching from target processes, we employ a strategy called “Check if Busy”. Specifically, when process A uses MPI_GET to fetch data from process B, both A and B are marked as BUSY. During this period, no other process can fetch data from either A or B; instead, they will first fetch data from processes that are NOT BUSY. This strategy helps prevent multiple processes from simultaneously accessing the same process, thereby reducing communication congestion and improving overall parallel efficiency.
3 Results
Although the TPSCI program in this work is designed for “QC+HPC”, this study is an independent effort and does not yet incorporate QC-based bit string selection. Instead, we perform FCI calculations to provide widely applicable benchmark data. Consequently, the computational scaling follows $N_{\mathrm{FCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}$, which is higher than the $N_{\mathrm{FCI}}\cdot N_{\mathrm{occ}}\cdot N_{\mathrm{vir}}$ scaling of the Handy-Knowles FCI algorithm. Our goal is not to compete with existing FCI methods but to showcase the effectiveness of the newly designed TPSCI algorithm and its high efficiency in parallel execution.
In actual calculations, we further introduced molecular symmetry to reduce both memory usage and computational cost. All HF reference states and their corresponding FCIDUMP files were generated using PySCF. 19 For double-precision TPSCI, convergence was defined as an energy difference of less than $1\times 10^{-7}$ Hartree between consecutive iterations. For single-precision TPSCI, due to fewer significant digits, a looser criterion of less than $1\times 10^{-6}$ Hartree was used. Notably, the final total energies obtained from both versions differ only in the sixth decimal place, indicating that the single-precision SCI calculation is accurate to $0.01$ milliHartree (mH), which is sufficient for chemical applications.
We perform experiments on supercomputer Fugaku, which is a CPU only architecture based on the A64FX processor. Each node has 48 cores, and only 32 GiB of memory, making processing and memory distribution essential. The code was compiled with the Fujitsu compilers using the -Kfast,parallel,openmp,SVE options.
3.1 N2 with the aug-cc-pVDZ Basis Set
Using $D_{2h}$ symmetry with frozen core orbitals, the N2 molecule with the aug-cc-pVDZ basis set has $1.47\times 10^{11}$ determinants. We first evaluate the computational time of a single Davidson iteration using different numbers of nodes. As shown in Table 1, the program’s wall time, defined as the wall time of the slowest process, remains within 116% of the average wall time across all processes, demonstrating good MPI load balancing.
Table 1: Parallel performance of TPSCI for N2 computed with the aug-cc-pVDZ basis set. We report in seconds the average wall-time (Avg.), maximum wall-time (Max.), and delay time ($T_{\text{delay}}$, see Sec. 2.4.4). We also report the total node hours (NHs) used by the calculation.
Nodes Avg. Max. $T_{\text{delay}}$ NHs
480 2259 2484 115 331.2
960 1179 1292 7 344.5
1440 834 917 5 366.8
1920 655 740 14 394.7
2400 554 618 11 412.0
2880 483 552 28 441.6
A particularly notable result is that the total delay time ($T_{\text{delay}}$) does not increase significantly as the number of nodes grows, demonstrating strong parallel efficiency. Generally, increasing the number of nodes worsens communication congestion, which in turn increases MPI latency. In fact, before implementing the optimization strategies introduced in Section 2.4, the measured $T_{\text{delay}}$ values exhibited a significant growth trend with increasing nodes, sometimes even resulting in longer wall times despite using more nodes. However, after applying these strategies, $T_{\text{delay}}$ was significantly reduced, allowing the wall time to decrease consistently with increasing nodes. Although the total node-hours (NHs) still increase with node count (i.e. decreasing parallel efficiency), the substantial wall time reduction confirms the effectiveness of our parallelization strategies.
Table 2: Potential Energy Surface of N2 in Hartree computed with the aug-cc-pVDZ basis set for different bond lengths (R). We report the HF energy and the FCI energies computed with double (FP64) and single (FP32) precision.
$R$ (Bohr) HF FP64 FP32
2.118 -108.956105 -109.297336 -109.297335
2.4 -108.875323 -109.258928 -109.258927
2.7 -108.748038 -109.182160 -109.182160
3.0 -108.618929 -109.108790 -109.108789
3.6 -108.401280 -109.016333 -109.016335
4.2 -108.242916 -109.986038 -109.986044
The FCI energies of the potential energy surface of N2 are shown in Table 2. The double-precision and single-precision results are nearly identical, with the maximum difference being only 0.006 mH at $R=4.2$ Bohr. This data validates the accuracy and reliability of our single-precision implementation for TPSCI calculations.
3.2 CN with the cc-pVTZ Basis Set
As our program is capable of performing open-shell system calculations, we also provide benchmark data for CN, computed with C2v symmetry, frozen core orbitals, and a bond length of 1.16 Å. For the cc-pVDZ and cc-pVTZ basis sets, these calculations require $2.46\times 10^{8}$ and $4.86\times 10^{11}$ determinants, respectively. We emphasize here that naively storing (in single precision) even a single CI vector alone on each node would require about 1.8 TB of memory; however, using our distributed memory scheme, we can unlock the full memory of supercomputer Fugaku and perform such a large calculation. We also report the strong scaling performance of this calculation in Table 3.
Table 3: Parallel performance of TPSCI for CN computed with the cc-pVTZ basis set. We report in seconds the average wall-time (Avg.), maximum wall-time (Max.), and delay time ($T_{\text{delay}}$, see Sec. 2.4.4). We also report the total node hours (NHs) used by the calculation.
Nodes Avg. Max. $T_{\text{delay}}$ NHs
2400 4019 4400 401 2933
4800 2231 2534 490 3379
7200 1497 1681 242 3362
In Table 4, we present a comparison of the energy values computed by FCI with various orders of coupled cluster theory computed using NWChem 20. The FCI values represent an improvement of approximately 1-2 mH over CCSDT 21, 22, 23. For a system of this size, CCSDTQ 24, 25 provides an excellent approximation, with an error of only about 0.2 mH. In addition, as our calculations represent one of only a few systems that can be computed with a triple-$\zeta$ quality basis set by FCI, we also perform $2/3$ extrapolation of this data to the complete basis set limit (CBS). We extrapolate the HF energy ($E_{c}^{\text{HF}}$) of a given cardinality ($c$) as CBS($l$/$m$) $=E^{\text{HF}}_{m}+\frac{(E^{\text{HF}}_{m}-E^{\text{HF}}_{l})m^{-5}}{m^{-5}-l^{-5}}$ and the correlation energy ($E_{c}^{\text{Cor}}$) as CBS($l$/$m$) $=E^{Cor}_{m}+\frac{(E^{\text{Cor}}_{m}-E^{\text{Cor}}_{l})m^{-3}}{m^{-3}-l^{-3}}$ 26. We hope that this data for a challenging system can be used for benchmarking new methods.
Table 4: Energy values of CN using different methods and basis sets. We denote the two point extrapolation using basis sets of cardinality $l$ and $m$ as CBS($l$/$m$).
Basis Set HF CCSD CCSD(T) CCSDT CCSDTQ FCI
cc-pVDZ -92.197616 -92.475781 -92.489728 -92.490269 -92.491601 -92.491812
cc-pVTZ -92.219490 -92.547372 -92.567450 -92.567478 -92.569054 -92.569252
CBS(2/3) -92.222807 -92.571623 -92.594283 -92.594095 -92.592456 -92.595966
4 Conclusion
Inspired by the idea of TPSCI proposed by an IBM team in 2024, we proposed and implemented a novel and efficient algorithm in this work. To overcome the memory limitations of a single node, we have, for the first time worldwide, implemented distributed storage of the CI vector in an SCI program, allowing the entire memory resources of a supercomputer to be utilized for large-scale computations. Additionally, to further reduce memory usage, we store only the diagonal elements of the Hamiltonian matrix while computing the off-diagonal elements on-the-fly. As a result, we successfully handled up to $4.86\times 10^{11}$ determinants, surpassing the previously reported largest SCI calculation ($2\times 10^{9}$ determinants) by more than two orders of magnitude.
Meanwhile, large-scale parallel computations on supercomputers face a critical challenge: increasing the number of nodes often leads to severe communication congestion, significantly degrading parallel efficiency. To address this, we have developed and implemented a series of optimization strategies tailored to TPSCI, including eliminating unnecessary communication, reducing long-distance data transfers, maintaining balanced computational load and memory distribution in MPI, hiding communication by prefetching, and detecting target process statuses before data transfer. The results demonstrate that even with up to 7,200 nodes, the program’s wall time remains within 116% of the average across all processes, and communication delays ($T_{\text{delay}}$) do not increase significantly with the number of nodes. Such exceptional performance lays a solid foundation for the development of TPSCI in the future.
Since this study is an independent effort of our research group, it does not yet incorporate determinant selection using QC. Therefore, we directly performed FCI calculations on N2 with the aug-cc-pVDZ basis set and CN with the cc-pVTZ basis set, both of which are larger than the cc-pVDZ basis set reported in previous FCI studies of these molecules. These data can be used as benchmark in future quantum chemistry studies.
Looking ahead, the efficient distributed and parallelized framework developed in this work provides a strong foundation for integrating TPSCI with QC approaches. In particular, quantum computers could be leveraged to identify important bit strings, which could then be used to construct a spin-adapted determinant space via tensor products. These determinants would be efficiently processed using our highly efficient and scalable program. The combination of quantum bit string selection with large-scale classical computation has the potential to greatly enhance electronic structure calculations, enabling the accurate simulation of even larger and more complex molecular systems.
Acknowledgments
The authors thank Muneaki Kamiya for suggesting a restart strategy in the Davidson algorithm to reduce memory usage, and Himari Pathak for discussions on applications.

### v2

@@
-Distributed implementation of tensor-product selected configuration interaction
+The First Selected Configuration Interaction Implementation with Distributed Storage of the Configuration Interaction Vector
 Abstract
-In recent years, the hybrid “QC+HPC” strategy—where quantum computers screen important determinants, followed by exact diagonalization on classical computers—has shown great potential in the study of strongly correlated systems in quantum chemistry.
-Last year, an IBM team proposed a novel scheme that utilizes quantum computers to select important bit strings that are then used to construct a spin-adapted determinant space via tensor products.
-Inspired by this, we have specifically designed a completely new algorithm for this tensor-product selected configuration interaction (SCI).
-Moreover, for the first time worldwide, we have implemented distributed storage of the CI vector in an SCI program, enabling efficient handling of large-scale computation.
-Since this study is independent and does not involve determinant selection by quantum computers, we employed our SCI program to conduct full configuration interaction (FCI) computations.
-Our FCI calculations for N2 (aug-cc-pVDZ) under D2h symmetry and CN (cc-pVTZ) under C2v symmetry, involving $1.47\times 10^{11}$ and $4.86\times 10^{11}$ determinants, respectively, exceed the previous record of $2\times 10^{9}$ determinants computed with the DICE program [J.
-Chem.
-Phys.
-149, 214110 (2018)] by more than two orders of magnitude.
-These results set a new benchmark for SCI computations and lay the groundwork for further advancements in SCI methods.
+[…]
 1 Introduction
-The Full Configuration Interaction (FCI) is an exact method for solving the Schrödinger equation in quantum chemistry.
-1, 2, 3, 4 However, due to its exponential scaling with system size, FCI remains limited to small molecules with small basis sets.
-By the use of a distributed CI vector, the largest system calculated to date is propane with the STO-3G basis set, involving up to $1.31\times 10^{13}$ determinants 4.
-In practical applications, the coupled cluster singles and doubles with perturbative triples [CCSD(T)] method 5 is known as the gold standard for weakly correlated systems owing to its balanced accuracy and efficiency.
-While achieving accurate calculations for strongly correlated systems remains the most formidable challenge in quantum chemistry, the selected configuration interaction (SCI) method has long been employed and is recognized as one of the most promising methods for addressing this issue.
-As early as 1973, Huron et al.
-introduced the Configuration Interaction using a Perturbative Selection made Iteratively (CIPSI) method 6 in which candidate determinants are selected by rigorously estimating their energy contributions via second-order perturbation theory.
-Although this selection scheme is considered reliable, its high computational cost precludes its application to larger molecules.
-Subsequently, methods such as Adaptive Sampling CI (ASCI) 7, 8, and Semistochastic Heat-Bath CI (SHCI) 9, 10, 11 were developed, which can essentially be regarded as approximations to CIPSI.
-In recent years, with the rise of quantum computing (QC), several research groups have attempted to rapidly screen important determinants using quantum devices.
-However, the noise inherent in quantum computers can lead to the loss of some determinants, preventing the generation of wavefunctions that strictly correspond to the eigenstates of the $\hat{S}^{2}$ operator.
-To address this issue, last year a team from IBM proposed an innovative strategy: using quantum computers to select important bit strings, and then constructing the complete determinants via tensor products.
-12 In this paper, we refer to this variant of SCI,which is characterized by this structure, as Tensor-Product SCI (TPSCI).
-Since CI calculations are variational, including more determinants generally improves accuracy; however, the associated memory and computational costs for diagonalization increase dramatically.
-Currently, the widely used DICE program, 9, 10 which distributes the nonzero Hamiltonian matrix elements across nodes, enables calculations with up to $2\times 10^{9}$ determinants.
-Nevertheless, DICE employs replicated storage for the CI vector, so that the memory available on a single node becomes a bottleneck, thereby limiting the maximum feasible CI vector size despite multi-node execution.
+Full Configuration Interaction (FCI) is an exact method for solving the Schrödinger equation in quantum chemistry.
+1, 2, 3, 4, 5 However, due to its exponential scaling with system size, FCI remains limited to small molecules with small basis sets.
+Enabled by distributed storage of the configuration interaction (CI) vector 2, as of 2024 FCI implementations have reached a scale of $1.31\times 10^{12}$ determinants.
+5
+In practical applications, the coupled cluster singles and doubles with perturbative triples [CCSD(T)] method 6 is widely regarded as the gold standard for weakly correlated systems, owing to its balanced accuracy and efficiency.
+In contrast, accurately describing strongly correlated systems remains one of the most formidable challenges in quantum chemistry.
+The selected configuration interaction (SCI) method has long been employed and is recognized as one of the most promising methods for addressing this issue.
+Over the past decades, several SCI methods have been proposed and developed, such as the Configuration Interaction using a Perturbative Selection made Iteratively (CIPSI), 7 Adaptive Sampling CI (ASCI), 8, 9 Semistochastic Heat-Bath CI (SHCI), 10, 11, 12 and iterative CI (iCI).
+13, 14, 15 While distributed storage of the CI vector has been successfully applied in FCI calculations, 2, 3, 5 its implementation in SCI poses significant challenges due to the irregular and dynamically evolving nature of the determinant space.
+To date, all existing SCI implementations have employed replicated CI vectors, which imposes a strict memory limit on the maximum feasible CI vector size, even when multiple nodes are used.
+As a consequence, the largest SCI calculation reported to date contains just $2\times 10^{9}$ determinants, far smaller than the $1.31\times 10^{12}$ determinants achieved in state-of-the-art FCI computations.
+Meanwhile, rapid progress in quantum computing (QC) has led several research groups to explore hybrid approaches that combine quantum and classical computers for studying strongly correlated systems.
+In 2023, QunaSys first proposed a strategy that uses quantum computers to screen important Slater determinants, followed by exact diagonalization on classical computers—a method referred to as Quantum-Selected Configuration Interaction (QSCI).
+16 A downside to this approach is that the noise in quantum computers may cause some determinants to be missed, making the wavefunction deviate from the eigenstates of the $\hat{S}^{2}$ operator.
+To overcome this issue, in 2024 an IBM team proposed an extension of the QSCI strategy in which quantum computers are used to select important bitstrings, after which the complete determinants are constructed via tensor products of the selected bitstrings.
+17 In this paper, we refer to this approach as Tensor-Product Bitstring SCI (TBSCI), as it is based on the tensor product structure of $\alpha$ and $\beta$ bitstrings.
 In this work, we have, for the first time worldwide, implemented distributed storage of the CI vector in an SCI program, overcoming single-node memory limitation and enabling the use of the entire supercomputer’s available memory for larger-scale calculations.
-Moreover, we have designed a completely new, highly efficient algorithm specifically for TPSCI, leveraging its unique tensor-product structure, and further optimized the program for efficient parallel execution on the supercomputer Fugaku.
-We also validate the use of single precision to reduce computational cost and memory use; this is a strategy pursued for Configuration Interaction as early as the work of Pakiari and Handy, 13 and in line with current trends in correlated methods.
+Moreover, we have designed an entirely new, highly efficient algorithm specifically for TBSCI, leveraging its unique tensor-product structure, and further optimized the program for efficient parallel execution on the supercomputer Fugaku.
+We also validate the use of single precision to reduce computational cost and memory use; this is a strategy pursued for Configuration Interaction as early as the work of Pakiari and Handy, 18 and in line with current trends in correlated methods.
 19 The combination of low memory footprint and low precision operations make our work promising for application to leadership class supercomputers based on GPUs.
-Ultimately, we computed the FCI energies of N2 (aug-cc-pVDZ) and CN (cc-pVTZ), which, under D2h and C2v symmetry, respectively, involve $1.47\times 10^{11}$ and $4.86\times 10^{11}$ determinants.
-These results significantly exceed the previously reported maximum SCI scale of $2\times 10^{9}$ determinants computed with the DICE program.
-11 These achievements lay a solid foundation for the future development and application of TPSCI.
-Meanwhile, as previous FCI calculations for N2 15 and CN 16 were performed with the smaller cc-pVDZ basis set, our results serve as valuable benchmark data for further research.
-2 CI Vector Distributed Algorithm for TPSCI
-In this section, we will describe the parallel algorithm we have designed for TPSCI, including the important optimizations required for large-scale application.
-2.1 Distributed CI Vector Storage
-In the framework proposed by IBM, only important bit strings are selected, which means that the chosen $\alpha$ and $\beta$ bit strings are identical.
-The determinant space constructed via tensor products can thus be used to generate closed-shell singlet wavefunctions that strictly correspond to the eigenstates of the $\hat{S}^{2}$ operator.
-In our program, the selected $\alpha$ and $\beta$ bit strings are allowed to differ.
-Suppose their sets are defined as $\{S_{w}^{\alpha}\mid w=1,2,\dots,x\}$ and $\{S_{u}^{\beta}\mid u=1,2,\dots,y\}$, where $S^{\alpha}$ and $S^{\beta}$ are the sets of selected $\alpha$-spin and $\beta$-spin bit strings, respectively, with elements denoted as $S_{w}^{\alpha}$ and $S_{u}^{\beta}$.
+Ultimately, we have used our implementation to compute: the potential energy surfaces (PESs) of N2 (aug-cc-pVDZ) and CN (cc-pVTZ), as well as single-point energies of Cr2 (STO-3G) and N2 (cc-pVTZ) near their equilibrium geometries.
+Among these systems, the largest case is N2 (cc-pVTZ), involving $2.62\times 10^{12}$ determinants under D2h symmetry with the frozen-core approximation.
+This scale surpasses the previous SCI record of $2\times 10^{9}$ determinants 12 by more than three orders of magnitude, and even doubles the size of the largest FCI calculation reported to date ($1.31\times 10^{12}$ determinants).
+5
+At present, our work focuses on implementing distributed storage of the CI vector in the Davidson diagonalization step of SCI——a particularly challenging component in the development of large-scale SCI methods.
+The subsequent second-order perturbative correction on top of SCI has not yet been implemented in such a distributed manner.
+As a result, our current program cannot perform a complete SCI calculation.
+Instead, we demonstrate its capabilities by performing FCI calculations, the results of which can serve as valuable benchmark data for future research in quantum chemistry.
+2 A Novel and Highly Efficient Algorithm for TBSCI
+In this section, we describe a novel and highly efficient algorithm specifically designed for TBSCI, which combines distributed storage of the CI vector with a suite of MPI optimization strategies to enable trillion-scale determinant calculations on modern supercomputers.
+2.1 Distributed Storage of CI Vector
+In the general framework of TBSCI, the sets of the selected bitstrings can be defined as $\{S_{w}^{\alpha}\mid w=1,2,\dots,l_{\alpha}\}$ and $\{S_{u}^{\beta}\mid u=1,2,\dots,l_{\beta}\}$, where $\{S^{\alpha}\}$ and $\{S^{\beta}\}$ are the sets of selected $\alpha$-spin and $\beta$-spin bitstrings, respectively, with elements denoted as $S_{w}^{\alpha}$ and $S_{u}^{\beta}$.
+Here, $l_{\alpha}$ denotes the number of selected $\alpha$ and $l_{\beta}$ the number of selected $\beta$ bitstrings.
 The determinant space constructed via tensor products is then given by:
-$\{|S_{w}^{\alpha}\rangle\otimes|S_{u}^{\beta}\rangle\mid w=1,\dots,x;\,u=1,\dots,y\}.$ (1)
+$\{|S_{w}^{\alpha}\rangle\otimes|S_{u}^{\beta}\rangle\mid w=1,\dots,l_{\alpha};\,u=1,\dots,l_{\beta}\}.$ (1)
 The ordering of $\{S^{\alpha}\}$ and $\{S^{\beta}\}$ can be freely chosen.
@@
 Specifically, the $K$-th determinant is denoted as $|D_{K}\rangle$ and given by:
-$|D_{K}\rangle=|S_{w}^{\alpha}\rangle\otimes|S_{u}^{\beta}\rangle,\quad w=\frac{K}{y}+1,\quad u=\text{MOD}(K,y).$ (2)
+$|D_{K}\rangle=|S_{w}^{\alpha}\rangle\otimes|S_{u}^{\beta}\rangle,\quad w=\frac{K}{l_{\beta}}+1,\quad u=\text{MOD}(K,l_{\beta}).$ (2)
+For convenience, $|D_{K}\rangle$ will be denoted as $|D_{(w,u)}\rangle$ throughout this work.
 When the number of determinants exceeds the storage capacity of a single node, the CI vector must be distributed across nodes.
@@
 In this work, we denote the total number of processes (each node runs exactly one process) as $N$.
-Let $m_{p}$ denote the number of $\alpha$ bit strings assigned to process $p$ (where $p=0,1,\dots,N-1$), with $\sum^{N-1}_{p=0}m_{p}=x$.
-Then, each process contains $m_{p}$ segments and thus is responsible for a total of $m_{p}\times y$ determinants.
+Let $m_{p}$ denote the number of $\alpha$ bitstrings assigned to process $p$ (where $p=0,1,\dots,N-1$), with $\sum^{N-1}_{p=0}m_{p}=l_{\alpha}$.
+Then, each process contains $m_{p}$ segments and thus is responsible for a total of $m_{p}\times l_{\beta}$ determinants.
 As each determinant is fully defined by its index through Eq.
@@
 Figure 1: Distribution of the CI vector into segments across $N$ processes.
-Each process holds a strict subset of $m_{p}$ of the $\alpha$ strings.
+Each process holds a strict subset of $m_{p}$ $\alpha$ strings.
 On a local process, the determinants are divided into segments (represented by vertical arrows), where a segment is composed of all determinants that share the same $\alpha$ string.
@@
 2.2 Distributed Matrix-Vector Multiplication
-In this work, we employ the traditional Davidson diagonalization method 17 for CI calculations, where the core computational step is evaluating $\mathbf{W}=H\cdot\mathbf{U}$, with Hamiltonian matrix $H$, and CI vectors $\mathbf{W}$ and $\mathbf{U}$.
-The most advanced SCI implementations to date follow two main strategies: some store the nonzero Hamiltonian elements in a distributed manner across nodes (e.g., DICE), while others compute the Hamiltonian matrix elements on-the-fly (e.g., QUANTUM 18).
+In this work, we employ the traditional Davidson diagonalization method 20 for CI calculations, where the core computational step is evaluating $\mathbf{W}=H\cdot\mathbf{U}$, with Hamiltonian matrix $H$, and CI vectors $\mathbf{W}$ and $\mathbf{U}$.
+The most advanced SCI implementations to date follow two main strategies: some store the nonzero Hamiltonian elements in a distributed manner across nodes (e.g., DICE 12), while others compute the Hamiltonian matrix elements on-the-fly (e.g., QUANTUM 21).
 However, both approaches store the CI vector entirely on a single node without exception.
-The advantage of these approaches is that each computing node performs a simple matrix-vector multiplication; at the end of each Davidson iteration, the effective Hamiltonian matrix is efficiently constructed using an MPI_ALLREDUCE operation.
+The advantage of these approaches is that each computing node performs a simple matrix-vector multiplication; at the end of each Davidson iteration, the effective Hamiltonian matrix can be efficiently constructed using an MPI_ALLREDUCE operation.
 However, a key limitation is that the CI vector size is strictly constrained by the memory of a single node, posing a major bottleneck in utilizing supercomputing resources for larger determinant spaces.
-In our program, the CI vector is distributed across nodes, and each node runs exactly one process.
+In our program, the CI vector is distributed across MPI processes, with exactly one process running on each node.
 In this setting, each node $p$ only stores its corresponding CI vector segments.
@@
 Since there are $N$ computing nodes, each Davidson iteration consists of $N$ steps.
-Unlike the implementation in QUANTUM, we precompute and store the diagonal elements of the Hamiltonian matrix in memory, while the off-diagonal elements are computed on-the-fly during the Davidson iterations.
+Unlike the implementation in QUANTUM, we precompute and store the diagonal elements of the Hamiltonian matrix across processes in a distributed manner consistent with the CI vector, while the off-diagonal elements are computed on-the-fly during the Davidson iterations.
 During the $N$ computation steps within a single Davidson iteration, each node operates independently without global synchronization.
@@
 For data communication, we adopt the MPI_GET operation, a type of one-sided communication, allowing nodes to fetch data asynchronously without interrupting the computation on the target node, further improving efficiency.
-2.3 Efficient Hamiltonian Computation for TPSCI
+2.3 Efficient Hamiltonian Computation for TBSCI
 Due to the Slater-Condon rules, $H$ is a highly sparse matrix, so minimizing operations that involve zero elements is crucial for efficiency.
@@
 Here, we first analyze the structure of Hamiltonian matrix elements in SCI, which provides the foundation for our efficient computation strategy.
-The Hamiltonian matrix element $\langle D_{w}|H|D_{u}\rangle$ can be expressed as $\langle S_{w}^{\alpha}|\langle S_{w}^{\beta}|H|S_{u}^{\alpha}\rangle|S_{u}^{\beta}\rangle$.
+Suppose the Hamiltonian matrix element we need to compute is $\langle D_{(w,u)}|H|D_{(x,v)}\rangle$, which can be expressed as $\langle S_{w}^{\alpha}|\langle S_{u}^{\beta}|H|S_{x}^{\alpha}\rangle|S_{v}^{\beta}\rangle$.
 According to the Slater-Condon rules, this matrix element can be nonzero only if:
-$\text{DIFF}(S_{w}^{\alpha},S_{u}^{\alpha})+\text{DIFF}(S_{w}^{\beta},S_{u}^{\beta})\leq 2,$ (4)
-where $\text{DIFF}(S_{w},S_{u})$ denotes the number of electron differences between bit strings $S_{w}$ and $S_{u}$.
-To classify these terms concisely, we introduce the notation $[a,b]$, where $a=\text{DIFF}(S_{w}^{\alpha},S_{u}^{\alpha})$ and $b=\text{DIFF}(S_{w}^{\beta},S_{u}^{\beta})$.
-For example, the term [1,1] corresponds to $\text{DIFF}(S_{w}^{\alpha},S_{u}^{\alpha})=1$ and $\text{DIFF}(S_{w}^{\beta},S_{u}^{\beta})=1$.
+$\text{DIFF}(S_{w}^{\alpha},S_{x}^{\alpha})+\text{DIFF}(S_{u}^{\beta},S_{v}^{\beta})\leq 2,$ (4)
+where $\text{DIFF}(S_{w},S_{x})$ denotes the number of electron differences between bitstrings $S_{w}$ and $S_{x}$.
+To classify these terms concisely, we introduce the notation $[a,b]$, where $a=\text{DIFF}(S_{w}^{\alpha},S_{x}^{\alpha})$ and $b=\text{DIFF}(S_{u}^{\beta},S_{v}^{\beta})$.
+For example, the term [1,1] corresponds to $\text{DIFF}(S_{w}^{\alpha},S_{x}^{\alpha})=1$ and $\text{DIFF}(S_{u}^{\beta},S_{v}^{\beta})=1$.
 Throughout this work, we use the shorthand notations [2,0], [1,1], [1,0], [0,2], [0,1], and [0,0] to represent a total of six terms in which the matrix elements may be nonzero.
@@
 1 for $S_{w}^{\alpha}$leftarrow $S^{\alpha}_{p,1}$ to $S^{\alpha}_{p,m_{p}}$ do
-2      for $S_{u}^{\alpha}$leftarrow $S^{\alpha}_{q,1}$ to $S^{\alpha}_{q,m_{q}}$ do
-3           case $\leftarrow$ DIFF($S_{w}^{\alpha}$, $S_{u}^{\alpha}$)
+2      for $S_{x}^{\alpha}$leftarrow $S^{\alpha}_{q,1}$ to $S^{\alpha}_{q,m_{q}}$ do
+3           case $\leftarrow$ DIFF($S_{w}^{\alpha}$, $S_{x}^{\alpha}$)
 4           if case = 2 then
-5                for $S_{w}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{y}$ do
-6                     $S_{u}^{\beta}\leftarrow S_{w}^{\beta}$
+5                for $S_{u}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{l_{\beta}}$ do
+6                     $S_{v}^{\beta}\leftarrow S_{u}^{\beta}$
 7                     Calculate [2,0] term
@@
 11                     else if case=1 then
-12                          for $S_{w}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{y}$ do
-13                               foreach $S_{u}^{\beta}$in SINGLE_LINK(:,$S_{w}^{\beta}$do
+12                          for $S_{u}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{l_{\beta}}$ do
+13                               foreach $S_{v}^{\beta}$in SINGLE_LINK(:,$S_{u}^{\beta}$do
 14                                    Calculate [1,1] term
@@
 16                                    end foreach
-17                                   $S_{u}^{\beta}\leftarrow S_{w}^{\beta}$
+17                                   $S_{v}^{\beta}\leftarrow S_{u}^{\beta}$
 18                                    Calculate [1,0] term
@@
 22                                    else if case=0 then
-23                                         for $S_{w}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{y}$ do
-24                                              foreach $S_{u}^{\beta}$in SINGLE_LINK(:,$S_{w}^{\beta}$do
+23                                         for $S_{u}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{l_{\beta}}$ do
+24                                              foreach $S_{v}^{\beta}$in SINGLE_LINK(:,$S_{u}^{\beta}$do
 25                                                   Calculate [0,1] term
@@
 27                                                   end foreach
-28                                                  $S_{u}^{\beta}\leftarrow S_{w}^{\beta}$
+28                                                  $S_{v}^{\beta}\leftarrow S_{u}^{\beta}$
 29                                                   Calculate [0,0] term
@@
 35                                                   end for
-36                                                  $S_{u}^{\alpha}$leftarrow $S_{w}^{\alpha}$
-37                                                   for $S_{w}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{y}$ do
-38                                                        foreach $S_{u}^{\beta}$in the q-th batch of DOUBLE_LINK(:,$S_{w}^{\beta}$do
+36                                                  $S_{x}^{\alpha}$leftarrow $S_{w}^{\alpha}$
+37                                                   foreach $S_{u}^{\beta}$in the $q$-th batch of {$S^{\beta}$do
+38                                                        foreach $S_{v}^{\beta}$in DOUBLE_LINK(:,$S_{u}^{\beta}$do
 39                                                             Calculate [0,2] term
@@
 42
-43                                                             end for
+43                                                             end foreach
 44
 45                                                             end for
-Algorithm 1 Computation algorithm for TPSCI on local process $p$ at step $q$.
-This algorithm describes how to compute $\langle D_{w}|H|D_{u}\rangle$, where $|D_{w}\rangle$ and $|D_{u}\rangle$ are determinants from processes $p$ and $q$, respectively.
-Our novel algorithm for TPSCI is shown in Algorithm 1.
+Algorithm 1 Conceptual algorithm for TBSCI on local process $p$ at step $q$.
+This algorithm describes how to compute $\langle S_{w}^{\alpha}|\langle S_{u}^{\beta}|H|S_{x}^{\alpha}\rangle|S_{v}^{\beta}\rangle$, where $\langle S_{w}^{\alpha}|\langle S_{u}^{\beta}|$ and $|S_{x}^{\alpha}\rangle|S_{v}^{\beta}\rangle$ are determinants from processes $p$ and $q$, respectively.
+For clarity, we present the algorithm without the symmetry optimization (Sec.
+2.4.2) and reassignment strategy (Sec.
+2.4.6).
+Our novel algorithm for TBSCI is shown in Algorithm 1.
 Each Davidson iteration consists of $N$ steps.
 In step $q$, the local process $p$ performs the computation using data fetched from process $q$, which is detailed in Algorithm 1.
-To eliminate unnecessary computations, we introduce the SINGLE_LINK and DOUBLE_LINK arrays, which store precomputed single and double excitation connections among $\{S_{w}^{\beta}\mid u=1,2,\dots,y\}$.
-These arrays ensure that Hamiltonian evaluations only occur when the electron difference between $|D_{w}\rangle$ and $|D_{u}\rangle$ does not exceed 2, directly eliminating invalid cases and improving efficiency.
-This optimization is enabled due to the unique structure of TPSCI: for each $\alpha$ bit string, the associated $\beta$ bit string set remains unchanged across determinants, allowing excitation connections to be precomputed and reused efficiently.
-As the DOUBLE_LINK array can be extremely large, it is distributed nearly evenly across all processes based on its indices.
-To keep track of this distribution, an auxiliary array is introduced to record the starting and ending indices of DOUBLE_LINK assigned to each process.
-Consequently, the computation of the [0,2] term is performed batch by batch over the $N$ steps (as shown in Lines 27-32 of Algorithm 1).
-Since the calculation of the [0,2] term is performed only in the loop over $S_{w}^{\alpha}$, each process $p$ will fetch data from the target process $q$ for the DOUBLE_LINK only once per Davidson step.
-Roughly speaking, the computational cost of TPSCI scales as $N_{\mathrm{SCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}$, where $N_{\mathrm{SCI}}$ is the number of selected determinants, $N_{\mathrm{occ}}$ the number of occupied orbitals, and $N_{\mathrm{vir}}$ the number of virtual orbitals.
-This can be further reduced by considering the sparsity of excitation links.
-The sparsity of single excitation links roughly scales as $\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}$, in which $N_{\mathrm{FCI}}$ is the FCI dimension, and thus the overall scaling is more precisely given by $N_{\mathrm{SCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot N_{\mathrm{SCI}}/N_{\mathrm{FCI}}$.
-2.4 Strategies for Efficient Parallel Execution in TPSCI
+To eliminate unnecessary computations, we introduce the SINGLE_LINK and DOUBLE_LINK arrays, which store precomputed single and double excitation connections among $\{S_{u}^{\beta}\mid u=1,2,\dots,l_{\beta}\}$.
+These arrays ensure that Hamiltonian evaluations only occur when the electron difference between $|D_{(w,u)}\rangle$ and $|D_{(x,v)}\rangle$ does not exceed 2, directly eliminating invalid cases and improving efficiency.
+This optimization is particularly effective due to the unique structure of TBSCI: for each $\alpha$ bitstring, the associated $\beta$ bitstring set remains unchanged across determinants, allowing excitation connections to be precomputed and reused efficiently.
+In practice, these connections are stored as one-dimensional arrays to minimize memory usage, accompanied by auxiliary index arrays that record the starting positions and lengths of each column.
+For ease of explanation, however, we describe them as two-dimensional arrays throughout this paper.
+As the DOUBLE_LINK array can be extremely large due to storing all double excitation connections between all $\beta$ bitstrings, it is divided into batches, one per process, each containing a subset of $\{S_{u}^{\beta}\mid u=1,2,\dots,l_{\beta}\}$.
+For each $S_{u}^{\beta}$, the corresponding column DOUBLE_LINK(:, $S_{u}^{\beta}$) is constructed and stored entirely on a single process, and an auxiliary array records the index range assigned to each process to facilitate data access.
+Consequently, the computation of the [0,2] term is carried out batch by batch over the $N$ steps (Lines 27–32 of Algorithm 1).
+Since this term involves only a loop over $S_{w}^{\alpha}$, without referencing $S_{x}^{\alpha}$, each process $p$ needs to fetch DOUBLE_LINK data from the target process $q$ only once per Davidson step.
+In contrast, the SINGLE_LINK array has to be replicated on each node; otherwise, MPI communication would be too frequent, as it is accessed over both $S_{w}^{\alpha}$ and $S_{x}^{\alpha}$ (Line 11 of Algorithm 1).
+It is reasonable to assume that the sparsities of single and double excitations scales as $\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}$.
+Thus, the storage requirement of SINGLE_LINK is approximately $l_{\beta}\cdot N_{\mathrm{occ}}\cdot N_{\mathrm{vir}}\cdot\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}$, where $N_{\mathrm{occ}}$ and $N_{\mathrm{vir}}$ denote the numbers of occupied and virtual orbitals, respectively.
+This may become a memory bottleneck when $l_{\beta}$ becomes large; in such a case, single excitations may be generated on-the-fly, at the cost of additional computation.
+However, since the limitation did not arise in the present study and may not occur in future applications—given that $\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}$ can be very small—we do not pursue its resolution here.
+As shown in Algorithm 1, this sparsity factor ($\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}$) appears once in the computations of the $[2,0]$ and $[0,2]$ terms, but twice in that of the $[1,1]$ term.
+Consequently, the dominant contributions come from the $[2,0]$ and $[0,2]$ terms, each scaling as $N_{\mathrm{SCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}$, while the $[1,1]$ term has a smaller scaling of $N_{\mathrm{SCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot N_{\mathrm{SCI}}/N_{\mathrm{FCI}}$.
+In practice, the $[0,2]$ term is the most time-consuming, as the information for double excitations between $S_{w}^{\beta}$ and $S_{u}^{\beta}$ is evaluated within the innermost loop.
+As mentioned earlier, our TBSCI program can also be used to perform FCI calculations.
+In this case, the sparsity factor becomes exactly one, and the computational cost simplifies to $N_{\mathrm{FCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}$, with the dominant contributions arising from the $[2,0]$, $[1,1]$, and $[0,2]$ terms.
+In practice, when our implementation is used for FCI the [1, 1] term is the most time-consuming.
+It should be emphasized that the computational cost of the Handy–Knowles FCI algorithm scales only as $N_{\mathrm{FCI}}\cdot N_{\mathrm{occ}}\cdot N_{\mathrm{vir}}$; as such, our intention is not to compete with existing FCI methods, but rather to showcase both the algorithmic effectiveness of the newly designed TBSCI method and the high efficiency enabled by our parallelization strategies.
+A summary of the scaling behavior and the dominant terms is provided in Table 1.
+Table 1: Computational scaling and dominant terms of SCI and FCI calculations in the framework of TBSCI.
+Method Overall scaling Dominant terms
+SCI $N_{\mathrm{SCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}$ $[2,0]$, $[0,2]$
+FCI $N_{\mathrm{FCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}$ $[2,0]$, $[1,1]$, $[0,2]$
+2.4 Strategies for Efficient Parallel Execution in TBSCI
 As described above, we distribute the CI vector across nodes to overcome single-node memory limitations, enabling computations to fully utilize the total available memory.
@@
 As the number of nodes increases, the overall MPI communication volume seemingly scales as $N_{\text{SCI}}^{2}$, making communication latency a potential bottleneck for computational efficiency.
-To address this issue, we have developed a series of strategies to reduce communication overhead and improve parallel execution.
+To address this issue, we have developed a suite of strategies to reduce communication overhead and improve parallel execution.
 2.4.1 Avoiding Unnecessary MPI Data Transfers
@@
 However, not all data from the target process is required.
-Specifically, if all $\alpha$ bit strings assigned to process $p$, namely $\{S^{\alpha}_{p,1},S^{\alpha}_{p,2},\dots,S^{\alpha}_{p,m_{p}}\}$, satisfy $\text{DIFF}(S^{\alpha}_{w},S^{\alpha}_{u})>2$ for any $S^{\alpha}_{w}$ in this set, then process $p$ does not need the segment corresponding to $S^{\alpha}_{u}$.
+Specifically, if all $\alpha$ bitstrings assigned to process $p$, namely $\{S^{\alpha}_{p,1},S^{\alpha}_{p,2},\dots,S^{\alpha}_{p,m_{p}}\}$, satisfy $\text{DIFF}(S^{\alpha}_{w},S^{\alpha}_{x})>2$ for any $S^{\alpha}_{w}$ in this set, then process $p$ does not need the segment corresponding to $S^{\alpha}_{x}$.
 In our implementation, such information is precomputed and stored, allowing each process to identify and skip unnecessary data transfers, thereby significantly improving communication efficiency.
-As a result, the MPI communication volume for process $p$ roughly scales as $m_{p}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}\cdot y$.
-It can be easily shown that the computational cost and the data communication volume on each process share the same scaling.
-Consequently, the overall MPI communication volume across all nodes scales as $N_{\mathrm{SCI}}$ rather than $N_{\mathrm{SCI}}^{2}$, providing a solid foundation for the efficient parallel execution of TPSCI on supercomputers.
+As a result, the MPI communication volume for process $p$ roughly scales as $m_{p}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}\cdot l_{\beta}$, matching the scaling of the local computational cost.
+Consequently, the total MPI communication volume across all nodes scales as $N_{\mathrm{SCI}}$, rather than $N_{\mathrm{SCI}}^{2}$, providing a solid foundation for the efficient parallel execution of TBSCI in large-scale calculations involving many nodes on supercomputers.
+2.4.2 Exploiting Molecular Symmetry
+Exploiting molecular symmetry is a well-established practice in quantum chemistry to enhance computational efficiency.
+Building upon the strategy of avoiding unnecessary MPI data transfers, leveraging molecular symmetry can further reduce the data transfer volume in the TBSCI framework.
+As shown in Algorithm 1, the segments satisfying $\text{DIFF}(S^{\alpha}_{w},S^{\alpha}_{x})=2$, which are typically much more than those satisfying $\text{DIFF}=1$, are used solely for computing the $[2,0]$ term.
+In this case, because the $\beta$ bitstrings must match exactly ($S^{\beta}_{v}\equiv S^{\beta}_{u}$), $S^{\alpha}_{w}$ and $S^{\alpha}_{x}$ must belong to the same symmetry sector, otherwise the corresponding Hamiltonian matrix element vanishes.
+Thus, molecular symmetry naturally restricts the possible communication targets and further reduces MPI communication.
+This strategy becomes particularly effective for molecules with high symmetry.
+For example, in systems with D2h symmetry, memory usage drops by $\sim$8×8$\times$8 ×, and computational cost and data transfer volume drop by $\sim$64×64$\times$64 ×.
+However, in this case, the selected determinants are no longer constructed as tensor products of independently selected $\alpha$ and $\beta$ bitstrings, but are further restricted—on top of the original tensor-product construction—to those consistent with the target symmetry.
+As a result, the storage layout no longer strictly follows the structure shown in Figure 1, and the computation is no longer performed strictly according to Algorithm 1, as only symmetry-allowed combinations are evaluated.
+Nevertheless, for clarity, we retain the original forms of both Figure 1 and Algorithm 1.
 2.4.3 Minimizing Long-Distance Communication in MPI
 Each node in the supercomputer is assigned a unique node identity number (ID) ranging from 00 to $N-1$.
-When suitably ordered, messages sent between nodes with closer IDs will on average require fewer hops compared to those with larger ID differences.
-Based on this, we sort the $\alpha$ bit strings according to their excitation levels relative to the Hartree-Fock (HF) $\alpha$ bit string, and then assign them to nodes in increasing order.
-Hence, lower-excitation $\alpha$ bit strings are assigned to lower-numbered nodes, whereas higher-excitation strings are assigned to higher-numbered nodes.
-As a result, communication between distant nodes is significantly reduced because the excitation differences between their assigned $\alpha$ bit strings are more likely to be all greater than 2, making communication unnecessary.
-This strategy is crucial for reducing MPI communication congestion, a well-known challenge in large-scale high-performance computing, by ensuring that each node primarily fetches data from nearby nodes, thereby improving parallel efficiency.
-2.4.3 Achieving Computational Load Balance in MPI
-In MPI-based parallel computing, maintaining computational load balance is crucial because the slowest process determines the overall execution time.
-The challenge is how to estimate each process’s workload in advance.
-As shown in Algorithm 1, the most time-consuming computation is the evaluation of the [1,1] term.
-Therefore, the number of single excitations associated with each $\alpha$ bit string can be used as a rough estimate of the computational workload for its corresponding segment.
-Consequently, the total computational workload of a process can be estimated by summing the single excitation counts of all $\alpha$ bit strings assigned to it.
-As described above, we assign the $\alpha$ bit strings to processes in increasing order of their excitation levels relative to the HF $\alpha$ bit string.
-This ordering is crucial in our TPSCI program to minimize long-distance communication between nodes.
-However, balancing memory usage and computational workload simultaneously presents a challenge.
-Since the total number of $\alpha$ bit strings is fixed, the average number per node is also predetermined.
-Similarly, the total number of single excitation counts of all $\alpha$ bit strings is fixed, meaning that the average single excitation count per process is also predetermined.
-To quantify the balance between memory and computation, we define two expansion factors: the memory expansion factor, which is the ratio of the assigned $\alpha$ bit string count to the average, and the computation expansion factor, which is the ratio of the assigned single excitation counts to the average.
-Enforcing strict memory balance (i.e., keeping the memory expansion factor close to 1) can lead to severe computational imbalance, causing some nodes to finish significantly earlier and remain idle while waiting for others.
-Conversely, enforcing strict computational balance (i.e., keeping the computation expansion factor close to 1) may result in a highly uneven memory distribution, with some nodes exceeding their memory limits, forcing an increase in the total number of nodes and leading to inefficient resource usage.
-To address this issue, we adopt a balanced strategy where both expansion factors are kept moderately greater than 1, ensuring that neither memory imbalance nor computational imbalance becomes excessive.
-2.4.4 Overlapping Computation and Data Transfer to Reduce MPI Latency
+When nodes are appropriately ordered, messages sent between nodes with closer IDs tend to require fewer network hops than those between nodes with larger ID differences.
+Leveraging this property, we sort the selected $\alpha$ bitstrings according to their excitation levels relative to the Hartree–Fock (HF) $\alpha$ bitstring, and assign them to nodes in increasing order.
+Hence, lower-excitation $\alpha$ bitstrings are allocated to lower-numbered nodes, while higher-excitation bitstrings are assigned to higher-numbered nodes.
+This excitation-aware mapping can significantly reduce MPI communication between distant nodes.
+Since the excitation differences between the $\alpha$ bitstrings assigned to distant nodes are more likely to exceed two, remote data retrieval is usually unnecessary.
+As a result, and quite remarkably, even in massively parallel jobs involving thousands or tens of thousands of nodes, each node primarily fetches data from its neighboring nodes.
+This locality-preserving strategy plays a critical role in alleviating MPI communication congestion and enables our SCI program to maintain high parallel efficiency even on computations involving thousands or tens of thousands of nodes.
+2.4.4 Achieving Memory and Computational Load Balance in MPI
+As shown in Algorithm 1, the computational cost associated with each segment corresponding to a given $S^{\alpha}_{w}$ can be readily estimated.
+Consequently, the total workload of a process can be approximated by summing the estimated costs of all $\alpha$ bitstrings assigned to it.
+As described above, we assign $\alpha$ bitstrings to processes in increasing order of their excitation levels relative to the HF $\alpha$ bitstring, a strategy that is crucial in our parallel execution.
+Based on this excitation-aware mapping, achieving simultaneous balance in memory usage and computational workload becomes a nontrivial problem.
+Since the total number of $\alpha$ bitstrings is fixed, the average number per process is predetermined.
+Likewise, the total estimated cost is fixed, implying that the average computational load per process is also predetermined.
+Therefore, we can define two expansion factors to quantify the tradeoff between memory and computation: the memory expansion factor, defined as the ratio of the assigned $\alpha$ bitstring count to the average, and the computation expansion factor, defined as the ratio of the estimated computational cost to the average.
+Enforcing strict memory balance (i.e., keeping the memory expansion factor close to 1) can result in severe computational imbalance, causing some processes to complete significantly earlier and remain idle while others continue.
+Conversely, enforcing strict computational balance (i.e., keeping the computation expansion factor close to 1) may lead to a highly uneven memory usage, with some nodes exceeding their memory limits, thereby forcing an increase in the total number of nodes and leading to inefficient resource usage.
+To address this issue, we adopt a balanced strategy where both expansion factors are maintained to be moderately greater than 1, ensuring that neither memory imbalance nor computational imbalance becomes excessive.
+In the special case of FCI calculations—used here solely for testing purposes—the computational cost per segment is exactly the same, and thus a simple uniform distribution strategy naturally ensures both memory and workload balance.
+2.4.5 Overlapping Computation and Data Transfer
 As mentioned above, each node in our program runs a single process.
@@
 Conversely, if $t_{\text{data}}\leq t_{\text{cal}}$, then $t_{\text{delay}}=0$, since MPI communication congestion during data transfer—if it occurs—does not prolong the wall time.
-The total delay time for a Davidson iteration, denoted as $T_{\text{delay}}$, is given by the sum of $t_{\text{delay}}$ over all $N$ steps.
+The total delay time $T_{\text{delay}}$ for a Davidson iteration is defined as the sum of $t_{\text{delay}}$ over all $N$ steps of a process.
+As the wall time is determined by the slowest process, we only report the $T_{\text{delay}}$ observed on that process in the Section 3.
+As the single-precision implementation involves only half the data volume of the double-precision case, its $T_{\text{delay}}$ is expected to be significantly smaller, as confirmed by our numerical experiments.
+2.4.6 Reassignment of $[0,2]$ for Delay Absorption
+Since data retrieval and computation are executed concurrently, delays ($t_{\text{delay}}>0$) are more likely to occur in steps where the computational time $t_{\text{cal}}$ is relatively small.
+Through careful analysis, we observe that the $[0,2]$ term (Lines 27–32 in Algorithm 1) exhibits a favorable computational pattern: while being one of the dominant terms in both SCI and FCI calculations, it incurs substantial computational cost but only negligible data retrieval.
+We therefore treat the $[0,2]$ tasks as a computational reservoir to absorb potential delays.
+Each process reassigns its $N$ batches of $[0,2]$ tasks to steps where $t_{\text{cal}}$ is pre-estimated to be small, thereby reducing both the probability and magnitude of $t_{\text{delay}}$.
+As discussed in Section 2.4.3, a local process may not need to retrieve any data from certain target processes due to the excitation-aware mapping.
+In such cases, delays are unlikely to occur, and the corresponding $[0,2]$ task for those steps are reassigned to other steps with small $t_{\text{cal}}$.
+Due to this reassignment performed prior to execution, the actual execution no longer strictly follows the “one step computes one batch of DOUBLE_LINK” structure outlined in Algorithm 1.
+Instead, some steps may process multiple batches of $[0,2]$ tasks, some may process none, and some may be entirely skipped by process $p$.
+Nevertheless, for clarity, we still retain the original form of Algorithm 1.
+2.4.7 Odd-Even Process Fetch Strategy
+A large number of simultaneous MPI data transfers can lead to communication peaks in the network.
+To alleviate this issue, we introduce an odd–even process fetch strategy.
+For each process, the target ranks are sorted by the number of required segments.
+Odd-numbered processes fetch data in ascending order, while even-numbered processes fetch in descending order.
+This simple strategy effectively reduces both the frequency and magnitude of $T_{\text{delay}}$, particularly in simulations involving thousands or tens of thousands of nodes.
 2.4.8 “Check if Busy” Strategy for Communication
@@
 This strategy helps prevent multiple processes from simultaneously accessing the same process, thereby reducing communication congestion and improving overall parallel efficiency.
+2.4.9 Sleep Strategy under Severe MPI Congestion
+Even with the above MPI optimization strategies, we still sometimes observe unexpectedly large values of $T_{\text{delay}}$ when using more than 5,000 nodes.
+Surprisingly, for the same system and input settings, $T_{\text{delay}}$ can be nearly zero in one run, but over 100 seconds in another.
+This instability is partly caused by external factors, such as MPI traffic from other jobs running at the same time, or different node placements by the job scheduler.
+However, we believe the deeper reason lies in the nature of MPI congestion.
+It is like a traffic jam: once congestion occurs, it induces blocking behavior, which in turn worsens further congestion.
+This kind of chain reaction becomes harder to control when using more nodes.
+To address this, we design a sleep strategy.
+Once a process finds that $t_{\text{delay}}>0.2$ seconds for a given step, it is interpreted as an indication of severe congestion.
+The process then sleeps for 0.1 seconds to mitigate further blocking.
+While sleeping, the process is marked as BUSY, so that other processes will preferentially fetch data from other sources.
+Naturally, the sleep time is included in the measured $T_{\text{delay}}$.
+Unlike the previous strategies, this sleep strategy may slightly increase the total $T_{\text{delay}}$ of a single Davidson iteration.
+However, in all of our tests, it consistently eliminates large delay events, making it a practical and effective safeguard.
+The specific thresholds of 0.2 and 0.1 seconds are chosen empirically based on our experience on Fugaku; different systems and applications may require fine-tuning of these parameters.
 3 Results
-Although the TPSCI program in this work is designed for “QC+HPC”, this study is an independent effort and does not yet incorporate QC-based bit string selection.
-Instead, we perform FCI calculations to provide widely applicable benchmark data.
-Consequently, the computational scaling follows $N_{\mathrm{FCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}$, which is higher than the $N_{\mathrm{FCI}}\cdot N_{\mathrm{occ}}\cdot N_{\mathrm{vir}}$ scaling of the Handy-Knowles FCI algorithm.
-Our goal is not to compete with existing FCI methods but to showcase the effectiveness of the newly designed TPSCI algorithm and its high efficiency in parallel execution.
-In actual calculations, we further introduced molecular symmetry to reduce both memory usage and computational cost.
-All HF reference states and their corresponding FCIDUMP files were generated using PySCF.
-19 For double-precision TPSCI, convergence was defined as an energy difference of less than $1\times 10^{-7}$ Hartree between consecutive iterations.
-For single-precision TPSCI, due to fewer significant digits, a looser criterion of less than $1\times 10^{-6}$ Hartree was used.
-Notably, the final total energies obtained from both versions differ only in the sixth decimal place, indicating that the single-precision SCI calculation is accurate to $0.01$ milliHartree (mH), which is sufficient for chemical applications.
-We perform experiments on supercomputer Fugaku, which is a CPU only architecture based on the A64FX processor.
-Each node has 48 cores, and only 32 GiB of memory, making processing and memory distribution essential.
-The code was compiled with the Fujitsu compilers using the -Kfast,parallel,openmp,SVE options.
+We first generated the HF reference states and corresponding FCIDUMP files using PySCF.
+22 Subsequently, TBSCI calculations were performed on the supercomputer Fugaku.
+The convergence criterion was defined as an energy difference of less than $2\times 10^{-7}$ Hartree between consecutive iterations.
+It should be noted that all the FCI data presented in this work were recomputed using our TBSCI program, although some of the results have been previously reported.
+23, 24 Fugaku features a CPU-only architecture based on the A64FX processor.
+Each node contains 48 cores and only 32 GiB of memory, making both computational and memory distribution essential for large-scale execution.
+The TBSCI code was compiled using the Fujitsu compiler with the -Kfast,parallel,openmp,SVE flags.
+To demonstrate the performance of our code, we report the time of a single distributed matrix-vector multiplication ($\mathbf{W}=H\cdot\mathbf{U}$) for different numbers of processes; the wall time of this multiplication is almost as same as that of one Davidson iteration.
+Throughout all calculations, all data, including the CI vector and intermediate arrays, were kept entirely in memory without disk I/O, ensuring minimal data access latency.
 3.1 N2 with the aug-cc-pVDZ Basis Set
-Using $D_{2h}$ symmetry with frozen core orbitals, the N2 molecule with the aug-cc-pVDZ basis set has $1.47\times 10^{11}$ determinants.
-We first evaluate the computational time of a single Davidson iteration using different numbers of nodes.
-As shown in Table 1, the program’s wall time, defined as the wall time of the slowest process, remains within 116% of the average wall time across all processes, demonstrating good MPI load balancing.
-Table 1: Parallel performance of TPSCI for N2 computed with the aug-cc-pVDZ basis set.
-We report in seconds the average wall-time (Avg.), maximum wall-time (Max.), and delay time ($T_{\text{delay}}$, see Sec.
-2.4.4).
-We also report the total node hours (NHs) used by the calculation.
+We first test the performance of TBSCI by computing the potential energy surface (PES) of N2 using the aug-cc-pVDZ basis set.
+Under D2h symmetry and the frozen-core approximation, each geometry point yields $1.47\times 10^{11}$ determinants.
+Table 2: Potential energy surface of N2 computed with the aug-cc-pVDZ basis set at various bond lengths ($R$).
+HF and FP64 (double-precision) FCI energies are reported in Hartree ($E_{h}$), and FP32 (single-precision) deviations from FP64 are shown in units of $10^{-6}\ E_{h}$.
+$R$ (Bohr) HF FP64 FP32
+2.068 -108.961045 -109.295263 1
+2.4 -108.875323 -109.258928 1
+2.7 -108.748038 -109.182160 0
+3.0 -108.618929 -109.108790 0
+3.6 -108.401280 -109.016333 -3
+4.2 -108.242916 -109.986039 -9
+[…]
+Table 3: Parallel performance of TBSCI for N2 computed with the aug-cc-pVDZ basis set using both double- and single-precision (FP64 and FP32) implementations.
+For each case, we report the average wall time (Avg.), maximum wall time (Max.), and delay time ($T_{\text{delay}}$, see Sec.
+2.4.5) of a single distributed matrix-vector multiplication in seconds, as well as the total node hours (NHs) consumed.
+Nodes FP64 FP32
+Avg. Max. $T_{\text{delay}}$ NHs Avg. Max. $T_{\text{delay}}$ NHs
+480 2030 2088 0 278 1942 1997 0 266
+960 1016 1053 0.1 281 968 1007 0 269
+1440 676 704 0.3 282 645 673 0 269
+1920 509 531 10.4 283 484 505 0 269
+2400 407 426 4.2 284 387 405 0 270
+2880 341 358 0.7 286 325 340 0 272
+The MPI performance of both precision variants is summarized in Table 3.
+We begin by analyzing the MPI communication delay, characterized by the total $T_{\text{delay}}$ on the slowest process.
+Both precision variants show satisfactory performance in this regard.
+Remarkably, $T_{\text{delay}}$ is exactly zero in all single-precision runs, highlighting the benefit of reduced data volume in mitigating MPI congestion.
+Next, we assess MPI load balancing using the ratio of maximum to average wall time.
+Across all node counts, this ratio remains within 105%, confirming that the uniform distribution of $\alpha$ bitstrings achieves both memory and computational balance in FCI calculations.
+Finally, we evaluate strong scaling efficiency by tracking the total node hours (NHs) with increasing node counts.
+As shown in Table 3, NHs increase only slightly, indicating that the implementation maintains near-ideal strong scaling up to almost 3,000 nodes.
+3.2 CN with the cc-pVTZ Basis Set
+The next tested system is an open-shell molecule, CN.
+Under C2v symmetry with frozen core orbitals, the number of determinants reaches $2.46\times 10^{8}$ and $4.86\times 10^{11}$ under the cc-pVDZ and cc-pVTZ basis sets, respectively.
+Table 4: Potential energy surfaces of CN computed with the cc-pVDZ and cc-pVTZ basis sets.
+HF and FP64 FCI energies are reported in Hartree ($E_{h}$), and FP32 deviations from FP64 are shown in units of $10^{-6}\ E_{h}$.
+The FP64 energies are extrapolated to the complete basis set (CBS) limit using a two-point formula, denoted as CBS($l$/$m$), where $l$ and $m$ denote the cardinalities of the basis sets.
+$R$ (Å) cc-pVDZ cc-pVTZ CBS(2/3)
+HF FP64 FP32 HF FP64 FP32 FP64
+1.16 -92.197616 -92.491811 0 -92.219490 -92.569252 0 -92.595966
+1.3 -92.149983 -92.479361 0 -92.168341 -92.550251 1 -92.575154
+1.5 -92.041711 -92.408658 1 -92.059594 -92.476056 1 -92.499617
+1.7 -91.937733 -92.334641 0 -91.957354 -92.400739 -1 -92.423284
+2.0 -91.822293 -92.262054 -1 -91.844595 -92.324826 -1 -92.345249
+2.3 -91.704156 -92.242169 4 -91.727232 -92.300497 -1 -92.318840
+[…]
+Table 5: Parallel performance of TBSCI for CN computed with the cc-pVTZ basis set using both FP64 and FP32 implementations.
+For each case, we report the average wall time (Avg.), maximum wall time (Max.), and delay time ($T_{\text{delay}}$) of a single distributed matrix-vector multiplication in seconds, as well as the total node hours (NHs) consumed.
+Nodes FP64 FP32
+Avg. Max. $T_{\text{delay}}$ NHs Avg. Max. $T_{\text{delay}}$ NHs
+1600 5516 5743 0 2552 5026 5265 0 2340
+2400 3655 3809 0 2539 3331 3478 0 2319
+4800 1824 1909 0.5 2545 1650 1743 0 2324
+7200 1227 1315 18.1 2630 1110 1224 0 2448
+9600 938 1070 8.3 2853 839 1003 0 2675
+The MPI performance of TBSCI for CN with the cc-pVTZ basis set is summarized in Table 5.
+While the double-precision variant shows modest delays at high node counts (e.g., 18.1 seconds at 7200 nodes), the single-precision variant consistently shows zero delay on the slowest process in all tests—even at 9,600 nodes.
+Such excellent performance once again highlights the advantage of reduced data volume in large-scale high-performance computing (HPC) environments.
+At 9600 nodes, the ratio of maximum to average wall time increases to nearly 120%, indicating growing MPI load imbalance.
+Although in FCI calculations each segment should, in principle, have the same computational cost, such balance cannot be maintained once the workload is divided into multiple steps.
+Because perfect parallelism cannot be achieved within each step, increasing the number of steps—an inevitable consequence of using more nodes—leads to wall-time differences between processes.
+As a result, MPI load imbalance becomes more significant at larger node counts, leading to the observed increase in NHs.
+Nevertheless, from 1600 to 9600 nodes, the total NHs increase by only 14%, highlighting the exceptional performance enabled by the MPI optimization strategies we have designed.
+3.3 Cr2 with the STO-3G Basis Set
+Cr2 is a challenging molecule in quantum chemistry, exhibiting strong static and dynamic correlation even near its equilibrium geometry.
+With the minimal STO-3G basis set, the FCI dimension still reaches $9.14\times 10^{11}$ under D2h symmetry with 12 core orbitals frozen.
+Table 6: Energies of Cr2 computed with various methods using the STO-3G basis set at $R=1.5$.
+Methods Energy
+HF -2064.078900
+CCSD(T) -2064.781602
+CCSDT -2064.780060
+CCSDTQ -2064.802602
+FP64 -2064.808195
+FP32 -2064.808195
+[…]
+Table 7: Parallel performance of TBSCI for Cr2 computed with the STO-3G basis set using the single-precision implementation.
+We report the average wall time (Avg.), maximum wall time (Max.), and delay time ($T_{\text{delay}}$, see Sec.
+2.4.5) of a single distributed matrix-vector multiplication in seconds, as well as the total node hours (NHs) consumed.
 Nodes Avg. Max. $T_{\text{delay}}$ NHs
-480 2259 2484 115 331.2
-960 1179 1292 7 344.5
-1440 834 917 5 366.8
-1920 655 740 14 394.7
-2400 554 618 11 412.0
-2880 483 552 28 441.6
-A particularly notable result is that the total delay time ($T_{\text{delay}}$) does not increase significantly as the number of nodes grows, demonstrating strong parallel efficiency.
-Generally, increasing the number of nodes worsens communication congestion, which in turn increases MPI latency.
-In fact, before implementing the optimization strategies introduced in Section 2.4, the measured $T_{\text{delay}}$ values exhibited a significant growth trend with increasing nodes, sometimes even resulting in longer wall times despite using more nodes.
-However, after applying these strategies, $T_{\text{delay}}$ was significantly reduced, allowing the wall time to decrease consistently with increasing nodes.
-Although the total node-hours (NHs) still increase with node count (i.e.
-decreasing parallel efficiency), the substantial wall time reduction confirms the effectiveness of our parallelization strategies.
-Table 2: Potential Energy Surface of N2 in Hartree computed with the aug-cc-pVDZ basis set for different bond lengths (R).
-We report the HF energy and the FCI energies computed with double (FP64) and single (FP32) precision.
-$R$ (Bohr) HF FP64 FP32
-2.118 -108.956105 -109.297336 -109.297335
-2.4 -108.875323 -109.258928 -109.258927
-2.7 -108.748038 -109.182160 -109.182160
-3.0 -108.618929 -109.108790 -109.108789
-3.6 -108.401280 -109.016333 -109.016335
-4.2 -108.242916 -109.986038 -109.986044
-The FCI energies of the potential energy surface of N2 are shown in Table 2.
-The double-precision and single-precision results are nearly identical, with the maximum difference being only 0.006 mH at $R=4.2$ Bohr.
-This data validates the accuracy and reliability of our single-precision implementation for TPSCI calculations.
-3.2 CN with the cc-pVTZ Basis Set
-As our program is capable of performing open-shell system calculations, we also provide benchmark data for CN, computed with C2v symmetry, frozen core orbitals, and a bond length of 1.16 Å.
-For the cc-pVDZ and cc-pVTZ basis sets, these calculations require $2.46\times 10^{8}$ and $4.86\times 10^{11}$ determinants, respectively.
-We emphasize here that naively storing (in single precision) even a single CI vector alone on each node would require about 1.8 TB of memory; however, using our distributed memory scheme, we can unlock the full memory of supercomputer Fugaku and perform such a large calculation.
-We also report the strong scaling performance of this calculation in Table 3.
-Table 3: Parallel performance of TPSCI for CN computed with the cc-pVTZ basis set.
-We report in seconds the average wall-time (Avg.), maximum wall-time (Max.), and delay time ($T_{\text{delay}}$, see Sec.
-2.4.4).
-We also report the total node hours (NHs) used by the calculation.
+2880 1098 1119 0.4 895
+5760 551 567 2.2 907
+8640 373 383 7.4 919
+11520 281 293 8.1 938
+14400 227 242 17.3 968
+For large-scale systems, we recommend using the single-precision implementation, as it provides sufficient numerical accuracy while significantly improving MPI efficiency compared to the double-precision variant.
+Therefore, all subsequent tests focus exclusively on the single-precision variant.
+As summarized in Table 7, $T_{\text{delay}}$ remains below 18 seconds even when using more than 14,000 nodes.
+The ratio of maximum to average wall time remains within 107%, and the total NHs increase by less than 9% from 2,880 to 14,400 nodes.
+These results demonstrate not only near-ideal parallel efficiency but also excellent scalability: even as both the system size and node count grow substantially, our implementation consistently maintains exceptional performance.
+This underscores the strength and robustness of our MPI optimization strategies in handling large-scale computations.
+3.4 N2 with the cc-pVTZ Basis Set
+The final system is N2 with the cc-pVTZ basis set at $R=2.068$ Bohr.
+This calculation involves $2.62\times 10^{12}$ determinants under D2h symmetry with core orbitals frozen, which is approximately twice the size of the largest reported FCI calculation to date ($1.31\times 10^{12}$ determinants).
+Table 8: HF and FCI energies (single-precision) of N2 at $R=2.068$ Bohr computed with the cc-pVDZ and cc-pVTZ basis sets.
+The FCI energies are extrapolated to the complete basis set (CBS) limit using a two-point formula.
+cc-pVDZ cc-pVTZ CBS(2/3)
+HF FP32 HF FP32 FP32
+-108.954553 -109.276527 -108.984093 -109.375153 -109.408722
+[…]
+Table 9: Parallel performance of TBSCI for N2 computed with the cc-pVTZ basis set using single-precision implementations.
+We report the average wall time (Avg.), maximum wall time (Max.), and delay time ($T_{\text{delay}}$) of a single distributed matrix-vector multiplication in seconds, as well as the total node hours (NHs) consumed.
 Nodes Avg. Max. $T_{\text{delay}}$ NHs
-2400 4019 4400 401 2933
-4800 2231 2534 490 3379
-7200 1497 1681 242 3362
-In Table 4, we present a comparison of the energy values computed by FCI with various orders of coupled cluster theory computed using NWChem 20.
-The FCI values represent an improvement of approximately 1-2 mH over CCSDT 21, 22, 23.
-For a system of this size, CCSDTQ 24, 25 provides an excellent approximation, with an error of only about 0.2 mH.
-In addition, as our calculations represent one of only a few systems that can be computed with a triple-$\zeta$ quality basis set by FCI, we also perform $2/3$ extrapolation of this data to the complete basis set limit (CBS).
-We extrapolate the HF energy ($E_{c}^{\text{HF}}$) of a given cardinality ($c$) as CBS($l$/$m$) $=E^{\text{HF}}_{m}+\frac{(E^{\text{HF}}_{m}-E^{\text{HF}}_{l})m^{-5}}{m^{-5}-l^{-5}}$ and the correlation energy ($E_{c}^{\text{Cor}}$) as CBS($l$/$m$) $=E^{Cor}_{m}+\frac{(E^{\text{Cor}}_{m}-E^{\text{Cor}}_{l})m^{-3}}{m^{-3}-l^{-3}}$ 26.
-We hope that this data for a challenging system can be used for benchmarking new methods.
-Table 4: Energy values of CN using different methods and basis sets. We denote the two point extrapolation using basis sets of cardinality $l$ and $m$ as CBS($l$/$m$).
-Basis Set HF CCSD CCSD(T) CCSDT CCSDTQ FCI
-cc-pVDZ -92.197616 -92.475781 -92.489728 -92.490269 -92.491601 -92.491812
-cc-pVTZ -92.219490 -92.547372 -92.567450 -92.567478 -92.569054 -92.569252
-CBS(2/3) -92.222807 -92.571623 -92.594283 -92.594095 -92.592456 -92.595966
-4 Conclusion
-Inspired by the idea of TPSCI proposed by an IBM team in 2024, we proposed and implemented a novel and efficient algorithm in this work.
-To overcome the memory limitations of a single node, we have, for the first time worldwide, implemented distributed storage of the CI vector in an SCI program, allowing the entire memory resources of a supercomputer to be utilized for large-scale computations.
-Additionally, to further reduce memory usage, we store only the diagonal elements of the Hamiltonian matrix while computing the off-diagonal elements on-the-fly.
-As a result, we successfully handled up to $4.86\times 10^{11}$ determinants, surpassing the previously reported largest SCI calculation ($2\times 10^{9}$ determinants) by more than two orders of magnitude.
-Meanwhile, large-scale parallel computations on supercomputers face a critical challenge: increasing the number of nodes often leads to severe communication congestion, significantly degrading parallel efficiency.
-To address this, we have developed and implemented a series of optimization strategies tailored to TPSCI, including eliminating unnecessary communication, reducing long-distance data transfers, maintaining balanced computational load and memory distribution in MPI, hiding communication by prefetching, and detecting target process statuses before data transfer.
-The results demonstrate that even with up to 7,200 nodes, the program’s wall time remains within 116% of the average across all processes, and communication delays ($T_{\text{delay}}$) do not increase significantly with the number of nodes.
-Such exceptional performance lays a solid foundation for the development of TPSCI in the future.
-Since this study is an independent effort of our research group, it does not yet incorporate determinant selection using QC.
-Therefore, we directly performed FCI calculations on N2 with the aug-cc-pVDZ basis set and CN with the cc-pVTZ basis set, both of which are larger than the cc-pVDZ basis set reported in previous FCI studies of these molecules.
-These data can be used as benchmark in future quantum chemistry studies.
-Looking ahead, the efficient distributed and parallelized framework developed in this work provides a strong foundation for integrating TPSCI with QC approaches.
-In particular, quantum computers could be leveraged to identify important bit strings, which could then be used to construct a spin-adapted determinant space via tensor products.
-These determinants would be efficiently processed using our highly efficient and scalable program.
-The combination of quantum bit string selection with large-scale classical computation has the potential to greatly enhance electronic structure calculations, enabling the accurate simulation of even larger and more complex molecular systems.
+12000 2582 2924 0.9 9747
+18000 1726 2076 0.3 10380
+24000 1346 1630 2.4 10876
+30000 1102 1413 12.8 11775
+36000 917 1201 8.2 12010
+As summarized in Table 9, $T_{\text{delay}}$ remains below 13 seconds even with more than 30,000 nodes, demonstrating the remarkable communication efficiency achieved by our MPI optimization strategies.
+However, as the number of nodes increases, the ratio of maximum to average wall time becomes significantly larger, reaching 131% at 36,000 nodes.
+Meanwhile, the total NHs increase about 23% from 12,000 to 36,000 nodes.
+Considering the use of tens of thousands of nodes, such load imbalance is acceptable, and the overall parallel performance remains excellent.
+Further improvements to the TBSCI program are planned for future work.
+4 Conclusions and Prospects
+In this work, we have proposed and implemented a completely new and highly efficient algorithm for the tensor-product bitstring SCI (TBSCI), in which the selected determinants are constructed as tensor products of important bitstrings.
+17 To overcome the memory limitations of a single node, we have, for the first time worldwide, implemented distributed storage of the CI vector in the TBSCI program, allowing the entire memory resources of a supercomputer to be utilized for large-scale computations.
+To reduce MPI communication congestion, we developed a suite of strategies including avoiding unnecessary communication, leveraging molecular symmetry, reducing long-distance data transfers, balancing computational load and memory usage, overlapping communication and computation through prefetching, reassigning fast-fetch $[0,2]$ tasks, implementing staggered fetch strategies, detecting process status prior to data access, and applying a sleep strategy.
+These innovations were not developed all at once, but were progressively identified and validated through several months of intensive testing, profiling, and iteration.
+As these strategies were integrated, the communication delay of the slowest process ($T_{\text{delay}}$) was drastically reduced.
+The parallel capabilities of TBSCI enabled an SCI-based calculation of N2 using the cc-pVTZ basis set with $2.62\times 10^{12}$ determinants—more than three orders of magnitude beyond the previous SCI record ($2\times 10^{9}$ determinants) 12 and about twice the scale of the largest reported FCI calculation ($1.31\times 10^{12}$ determinants).
+5 Despite this scale, the TBSCI single-precision implementation exhibited no communication delay under 10,000 nodes, and maintained an exceptional performance with under 18 seconds delay even at 36,000 nodes, marking a new level of MPI efficiency in large-scale SCI.
+Looking forward, a critical next step is the incorporation of second-order perturbative corrections atop the distributed SCI framework.
+Because TBSCI is the first SCI implementation with distributed CI vector storage, corresponding perturbative corrections must also be implemented in a distributed manner, requiring new algorithms and implementations.
+This presents substantial technical challenges which will be the subject of future work.
+Additionally, we are exploring the development of a fully classical bitstring selection strategy, which would enable TBSCI to operate entirely on classical hardware.
+Such a version may even outperform hybrid quantum-classical approaches by avoiding the detrimental effects of quantum noise and enabling complete reproducibility.
+Finally, the distributed storage scheme and MPI communication optimization strategies developed in this work are not limited to TBSCI.
+Rather, they constitute a general and scalable framework for large-scale CI implementations, and can be readily adapted to various SCI variants, as well as semistochastic CI methods.
+33, 34
+We believe this work lays the foundation for a new generation of parallel electronic structure methods capable of addressing strongly correlated systems at unprecedented scale.
+It also marks a new stage for SCI, driven by distributed CI vector storage and communication optimization strategies.
 Acknowledgments
-The authors thank Muneaki Kamiya for suggesting a restart strategy in the Davidson algorithm to reduce memory usage, and Himari Pathak for discussions on applications.
+The authors thank Muneaki Kamiya for suggesting a subspace restart strategy in the Davidson algorithm to reduce memory usage, and Himadri Pathak for discussions on applications